OpenStudy (anonymous):

solve the equation: x-3-4√x=0

OpenStudy (anonymous):

x-3 = 4 root x squaring both sides - x2-6x+9 = 16x x2-22x+9 = 0 solve accordingly

OpenStudy (anonymous):

k